# Enhancing In-Context Learning Performance with just SVD-Based Weight Pruning: A Theoretical Perspective

### v1

@@ Abstract
Pre-trained large language models (LLMs) based on Transformer have demonstrated striking in-context learning (ICL) abilities. With a few demonstration input-label pairs, they can predict the label for an unseen input without any parameter updates. In this paper, we show an exciting phenomenon that SVD-based weight pruning can enhance ICL performance, and more surprising, pruning weights in deep layers often results in more stable performance improvements in shallow layers. However, the underlying mechanism of those findings still remains an open question. To reveal those findings, we conduct an in-depth theoretical analysis by presenting the implicit gradient descent (GD) trajectories of ICL and giving the mutual information based generalization bounds of ICL via full implicit GD trajectories. This helps us reasonably explain the surprising experimental findings. Besides, based on all our experimental and theoretical insights, we intuitively propose a simple, model-compression and derivative-free algorithm for downstream tasks in enhancing ICL inference. Experiments on benchmark datasets and open source LLMs display the method effectiveness [^1] The code is available at https://github.com/chen123CtrlS/EnhancingICL_SVDPruning.
@@ 1 Introduction
In this paper, we show our surprising findings in ICL inference by experimentally analyzing the effect of singular value decomposition (SVD)-based pruning on the model performance at different depth layers. As demonstrated in Figure 1: (i) SVD-based weight pruning can enhance ICL performance. Across all cases, it is evident that SVD-based weight pruning can generally enhance model performance at various depth layers, compared to the baseline performance without weight pruning (indicated by the dashed lines); (ii) Pruning weights in deep layers often results in more stable performance improvements in shallow layers. Specially, deep layers weight matrices can be drastically reduced without much degradation in model performance, and may even get large improvements on it, while the model performance collapses after a sharp reduction at the shallow layers (see Section 2 for details). A similar case in Sharma et al. [36] notes that a large portion of singular values can be removed from linear layers in large language models without affecting or even improving reasoning performance. However, the underlying mechanism of this phenomenon still remains a mystery, and this paper seeks to explore the issue from the following two aspects.
Figure 1: The effect of weight pruning across different layer types. The figure shows the phenomenon observed on the benchmark datasets (SST-2, RTE, COPA) and open source LLMs (GPT-J-6B and LLAMA2-7B). Each sub-figure corresponds only to the indicated type of dataset, model and module. Notice that this figure mainly focuses on exhibiting the impact of weight pruning to the first two and the last two layers of the model and different colors are used to distinguish between these layers. The dashed line represents the pretrained model performance without SVD. We operate on the whole of MLP or ATTN and specifically marked the points of highest performance. The amount of weight pruning is severe, for instance, the highest model performance sometimes occurs at a clipping rate of 0.995. This is about 99.5% of the matrix’s original rank.
@@ 1.1 Related Works
To be brief, our work is intimately related to the lines of work on low-rank properties of deep neural networks, in-context learning and gradient descent, as well as information-theoretic generalization bounds. Due to limited space, we discuss these related works in detail in Appendix A.
@@ 2 SVD-Based Weight Pruning can Enhance ICL
The optimal rank-$r$ approximation and SVD. Given a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ and a constant $r\leq\min(m,n),r\in\mathbb{N}$. Eckart–Young–Mirsky theorem [13] provides an optimal solution $\mathbf{W}^{*}(r)=\mathbf{U}_{r}\mathbf{\Sigma}_{r}\mathbf{V}_{r}^{T}$, $\mathbf{U}_{r},\mathbf{V}_{r}$ are matrices containing the singular vectors corresponding to the largest $r$ singular values $[\sigma]_{1}^{r}$. Let $\xi=1-\frac{r}{\min(m,n)}\in(0,1)$ be the clipping rate.
@@ 2.1 Effect of Weight Pruning across Different Layer and Module Types
All sub-figures clearly show an interesting phenomenon about these models in ICL inference: SVD-based weight pruning can enhance ICL performance in both shallow and deep layers across different module types. More surprising, deep layers weight matrices can be drastically reduced without much degradation in model performance, and may even get large improvements on it. This suggests that pruning weights in deep layers can effectively reduce model complexity while maintaining or enhancing model stability. And the model performance collapses to 0.5 (the expectation of a random guess in the binary task) after a sharp reduction at the shallow layers by contrast, indicating a higher sensitivity of the model to changes in shallow layers.
@@ 2.2 Effect of Different ICL Shot Numbers
Firstly, we note that without weight pruning, the performance of the model improves with an increase in the number of ICL shots. Which is consistent with prior works. Besides, for every shot number, a phenomenon consistent with what is described in Section 2.1 is observed: SVD-based weight pruning can enhance ICL performance, pruning deep layer weight is more stable than pruning shallow weight. Last but not the least, Figure 2 also demonstrates roughly that with a decrease in the number of shots, the rate of performance collapse in the model slows down after a sharp reduction at the shallow layer. Intuitively, this is because LLMs exhibit a shift in focus in the ICL setting, which results in a reduced scope of the output space. The more shots there are, the more pronounced the shift becomes, and this also leads to a faster collapse. We also theoretically discuss it in Section 3.2.
@@ 3 Theoretical Analysis Results
In this Section, we first describe the core components of our study by reviewing some basic notations in Section 3.1, and present the implicit gradient descent trajectories of ICL in Section 3.2. Afterwards, we give a mutual information based generalization bounds of ICL via full implicit GD trajectories in Section 3.3. Based on all our experimental and theoretical insights, we intuitively propose a derivative-free and effective method for downstream tasks in enhancing ICL inference in Section 3.4. Complete proofs can be found in the Appendix. For ease of qualitative analysis, our theoretical analysis is mainly focuses on linear attention setting. We discuss the standard Softmax attention setting in Appendix B.2 and feed-forward (MLP) layers in Appendix B.3.
@@ 3.1 Preliminaries
Expected generalization error. Given a loss function $\ell:\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}^{+}$, where $\ell(w,\mathbf{h})$ measures the "unfitness" or "error" of any $\mathbf{h}\in\mathcal{H}$ with respect to a hypothesis $w\in\mathcal{W}$. We take $\ell$ as a continuous function and assume that $\ell$ is differentiable almost everywhere with respect to $w$. The goal of learning is to find a hypothesis $w$ that minimizes the population risk, and for any $w\in\mathcal{W}$, the population risk is defined as $L_{\mu}(w)\triangleq\mathbb{E}_{\mathbf{h}\sim\mu}[\ell(w,\mathbf{h})].$ However, since only can partially observe $\mu$ via the sample $\mathbf{H}_{s}$, we instead turn to use the empirical risk, defined as $L_{\mathbf{H}_{s}}(w)\triangleq\frac{1}{N}\sum_{i=1}^{N}\ell(w,\mathbf{h}_{i})$. Then the expected generalization error of $\mathcal{A}$ is defined as
@@
In-context learning with Transformers. By prompt design, most context semantic understanding tasks can be unified into classification tasks. Simplify the form of each token in $\mathbf{H}$ to $\mathbf{h}_{i}=[\mathbf{x}_{i},\mathbf{y}_{i}]$, where $[\mathbf{x}_{i}]_{1}^{N}\in\mathbb{R}^{din}$ and $[\mathbf{y}_{i}]_{1}^{N}\in\mathbb{R}^{dout}$ are encoded input text and corresponding labels respectively. The test token has the form $\mathbf{h}_{N+1}=[\mathbf{x}_{N+1},mask]$, where $mask$ is the label needed to predict, and usually set 00 as its initialization. Therefore, the form of attention with residual connection222 In real-world scenarios, the residual connection module in Transformers is indispensable. is as follows:
@@ 3.2 The Implicit Gradient Descent Trajectories of ICL
To begin with, we present the implicit gradient descent of ICL, inspired by [19, 10, 33]. These works describe how ICL with attention can be connected to a meta-optimizer produces implicit gradient. The following lemma demonstrates the result.
@@
Consider a Transformer consists of a single linear layer attention with residual connection, parameterized by $\mathbf{W}_{V}$,$\mathbf{W}_{K}$,$\mathbf{W}_{Q}$ as in Eq.(3). Same to Section 3.1, let $\mathbf{H}=[\mathbf{H}_{s},\mathbf{h}_{N+1}]$ be the input, where $\mathbf{H}_{s}=[\mathbf{h}_{1},\mathbf{h}_{2},...,\mathbf{h}_{N}]$ represents the demonstration sample and $\mathbf{h}_{N+1}$ is the test data. And let $\hat{\mathbf{h}}_{N+1}$ be the single layer output. Then, it holds that
@@
where $\mathbf{W}_{V}\mathbf{H}_{s}$ is regarded as the meta-gradient of ICL, which is used to generate the implicit gradient matrix $\Delta\mathbf{W}_{icl}$ to act on the final feature output. See Appendix B.1 for a proof.
@@
It is worth noting that $\text{rank}(\Delta\mathbf{W}_{icl})\leq\text{rank}(\mathbf{H}_{s})\leq N$ by rank relations in matrix multiplication, indicating that this is a low-rank operation. This explains why the effect of different ICL shot numbers is not uniform in Section 2.2. Details of the discussion in standard Softmax attention setting can be found in Appendix B.2. Plus, a significant body of work has discovered that the ICL capabilities of models are not robust to the order of ICL sample. For instance, Lu et al. [26] observed that large language models (LLMs) are sensitive to the sequence of ICL examples, and Liu et al. [25] reported that the context-based ICL performance exhibits a U-shaped curve—models tend to perform better with information that appears at the beginning or at the end of the input context. This aligns with observations from actual model training where the sequence of training samples affects outcomes especially when training with a batch size of $1$.
@@
where $\Delta\mathbf{W}_{t}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}^{t}{\mathbf{h}_{i}^{t-1}}\otimes\mathbf{W}_{K}^{t}{\mathbf{h}_{i}^{t-1}}\right)\mathbf{W}_{Q}^{t},$ $\mathbf{W}_{0}=0,\mathbf{W}_{t}=\mathbf{W}_{t-1}+\mathbf{G}_{t}$ and $\mathbf{G_{1}}=\Delta\mathbf{W}_{1}$. See Appendix B.4 for a proof.
@@ 3.3 Generalization Bounds of ICL via Full Implicit GD Trajectories
In addition, we also make a discussion on the noise of the implicit gradient in Section C.4, our analytical discussion indicates that the implicit gradients produced by Transformers in practical applications are noisy due to factors such as the extent of model pre-training and data characteristics (eg. ICL shot number). We first present the assumption used in this subsection.
@@
where $\mathbf{vec(G}_{t})\in\mathbb{R}^{d}$, $\Delta\mathbf{W}_{t}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}^{t}{\mathbf{h}_{i}^{t-1}}\otimes\mathbf{W}_{K}^{t}{\mathbf{h}_{i}^{t-1}}\right)\mathbf{W}_{Q}^{t}$ and $\mathbf{G}_{t}=\Delta\mathbf{W}_{t}(1+\mathbf{W}_{0}+\sum_{j=1}^{t-1}\mathbf{G}_{j})=\Delta\mathbf{W}_{t}(1+\mathbf{W}_{t-1})$. $\mathrm{tr}\{\cdot\}$ denotes the trace of a matrix, $\left\lVert\cdot\right\rVert_{F}$ denotes the Frobenius norm of a matrix and $\mathbb{E}^{X}_{Y}$ is the conditional expectation. Proof details in Appendix B.5.
@@
Theorem 2 indicates that one can control the generalization performance of ICL via controlling the implicit gradient norm along the entire ICL implicit GD trajectories. Specifically, modulating the norm of $[\mathbf{G}_{t}]_{1}^{L}$ or $[\Delta\mathbf{W}_{t}]_{1}^{L}$ may enhance performance when utilizing ICL. Note that controlling implicit gradient norm can also control the magnitude of the trace of implicit gradient noise covariance. This elucidates why weight pruning through SVD, even if it only alters a single weight matrix, can confer advantages on the performance of Transformers in ICL inference. We will present an example below demonstrating how weight pruning can affect the norm of $\mathbf{G}_{t}$ or $\Delta\mathbf{W}_{t}$, thereby influencing the expected generalization error. Additional examples provided in Appendix B.6. This is also why, as illustrated in Figure 1, the highest model performance sometimes occurs at a clipping rate of 0.995. Furthermore, this could elucidate the utility of normalization in Transformers. However, weight pruning also impacts the expressive power, implying that increased pruning does not invariably lead to better outcomes. This clarifies why the highest model performance may also occur at clipping rates lower than 0.995.
@@
Consider $\Delta\mathbf{W}_{k}=\left(\sum_{i=1}^{N}\mathbf{W}_{V}^{k}{\mathbf{h}_{i}^{k-1}}\otimes\mathbf{W}_{K}^{k}{\mathbf{h}_{i}^{k-1}}\right)\mathbf{W}_{Q}^{k}$ when $t=k$. And we primarily consider the changes in a upper bound written as $UB(\left\lVert\Delta\mathbf{W}_{k}\right\rVert_{F}^{2})\triangleq\sum_{i=1}^{N}\left\lVert\mathbf{W}_{V}^{k}{\mathbf{h}_{i}^{k-1}}\otimes\mathbf{W}_{K}^{k}{\mathbf{h}_{i}^{k-1}}\right\rVert_{F}^{2}\left\lVert\mathbf{W}_{Q}^{k}\right\rVert_{F}^{2}$. Let $r$ represents the remained rank and $\delta$ represents the potential noise consisting of parts with small singular values, that is, $\mathbf{W}_{V}^{k}=\mathbf{W}_{V_{r}}^{k}+\delta_{V}=\mathbf{U}_{r}^{V}\mathbf{\Sigma}_{r}^{V}(\mathbf{V}_{r}^{V})^{T}+\delta_{V}$, the same operation is applied to $\mathbf{W}_{Q}^{k}$ and $\mathbf{W}_{K}^{k}$ as well. Then we have
$\left\lVert\mathbf{W}_{V}^{k}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}$ $=\left\lVert\mathbf{W}_{V_{r}}^{k}+\delta_{V}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}$
$=\left\lVert\mathbf{W}_{V_{r}}^{k}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}+\left\lVert\delta_{V}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}$
$\geq\left\lVert\mathbf{W}_{V_{r}}^{k}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2},$
and further have
@@
where Eq. (*) is by
@@
and Eq. (**) is by $\left\lVert\mathbf{A}\right\rVert_{F}^{2}=\sum_{i}\sigma_{i}(\mathbf{A})$ for any matrix $\mathbf{A,B}$. $UB(\left\lVert\Delta\mathbf{W}_{k}(r)\right\rVert_{F}^{2})$ is the upper bound on $\lVert\Delta\mathbf{W}_{k}\lVert_{F}^{2}$ after using SVD.

### v2

@@ Abstract
Pre-trained large language models (LLMs) based on Transformer have demonstrated striking in-context learning (ICL) abilities. With a few demonstration input-label pairs, they can predict the label for an unseen input without any parameter updates. In this paper, we show an exciting phenomenon that SVD-based weight pruning can enhance ICL performance, and more surprising, pruning weights in deep layers often results in more stable performance improvements than in shallow layers. However, the underlying mechanism of those findings still remains an open question. To reveal those findings, we conduct an in-depth theoretical analysis by presenting the implicit gradient descent (GD) trajectories of ICL and giving the mutual information based generalization bounds of ICL via full implicit GD trajectories. This helps us reasonably explain the surprising experimental findings. Besides, based on all our experimental and theoretical insights, we intuitively propose a simple, model-compression and derivative-free algorithm for downstream tasks in enhancing ICL inference. Experiments on benchmark datasets and open source LLMs display the method effectiveness [^1] The code is available at https://github.com/chen123CtrlS/EnhancingICL_SVDPruning..
@@ 1 Introduction
In this paper, we show our surprising findings in ICL inference by experimentally analyzing the effect of singular value decomposition (SVD)-based pruning on the model performance at different depth layers. As demonstrated in Figure 1: (i) SVD-based weight pruning can enhance ICL performance. Across all cases, it is evident that SVD-based weight pruning can generally enhance model performance at various depth layers, compared to the baseline performance without weight pruning (indicated by the dashed lines); (ii) Pruning weights in deep layers often results in more stable performance improvements than in shallow layers. Specially, deep layers weight matrices can be drastically reduced without much degradation in model performance, and may even get large improvements on it, while the model performance collapses after a sharp reduction at the shallow layers (see Section 2 for details). A similar case in Sharma et al. [38] notes that a large portion of singular values can be removed from linear layers in large language models without affecting or even improving reasoning performance. However, the underlying mechanism of this phenomenon still remains a mystery, and this paper seeks to explore the issue from the following two aspects.
Figure 1: The effect of weight pruning across different layer types. The figure shows the phenomenon observed on the benchmark datasets (SST-2, RTE, COPA) and open source LLMs (GPT-J-6B and LLAMA2-7B). Each sub-figure corresponds only to the indicated type of dataset, model and module. Notice that this figure mainly focuses on exhibiting the impact of weight pruning to the first two and the last two layers of the model and different colors are used to distinguish between these layers. The dashed line represents the pretrained model performance without SVD. We operate on the whole of MLP or ATTN and specifically marked the points of highest performance. The amount of weight pruning is severe, for instance, the highest model performance sometimes occurs at a clipping rate of 0.995. This is about 99.5% of the matrix’s original rank. For the definitions of “deep” and “shallow”, please refer to Appendix B.2.
@@
Our primary objective is to establish a general theoretical framework that uncovers the underlying mechanism behind the phenomenon that SVD-based weight pruning can enhance ICL performance. Based on our theoretical insights, one can design new ICL algorithms. Accordingly, we did not directly compare our approach with other pruning methods. The algorithm in Section 3.4 is presented solely to illustrate how theoretical analysis can guide experimental procedures effectively.
@@ 1.1 Related Works
Model compression. In recent years, there has been growing theoretical and experimental evidence that models can be significantly pruned with very little drop in accuracy, thereby significantly reducing their storage requirements. To name a few, Frankle and Carbin [15] indicate that neural networks can typically have over 90% of their weights eliminated with little to no loss in performance. Analogous results have been demonstrated in both feed-forward and convolutional networks used for image classification [24, 49, 22, 37]. More specifically, Sharma et al. [38] present that careful pruning done at specific layers of Transformer models can produce significant boosts in performance on some tasks. The discovery of this phenomenon heightens interest in the connection between generalization and over-parameterization [56, 57], it also spurs research into developing pruning strategies that facilitate efficient model inference [30]. Additionally, the works [19, 40] find that performance remains nearly unchanged until a significant portion (up to half) of the layers are removed.
@@ 2 SVD-Based Weight Pruning can Enhance ICL
The optimal rank-$r$ approximation and SVD. Given a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ and a constant $r\leq\min(m,n),r\in\mathbb{N}$. Eckart–Young–Mirsky theorem [14] provides an optimal solution $\mathbf{W}^{*}(r)=\mathbf{U}_{:r}\mathbf{\Sigma}_{:r}\mathbf{V}_{:r}^{T}$, $\mathbf{U}_{:r},\mathbf{V}_{:r}$ are matrices containing the singular vectors corresponding to the largest $r$ singular values $[\sigma]_{1}^{r}$. Let $\xi=1-\frac{r}{\min(m,n)}\in(0,1)$ be the clipping rate.
@@ 2.1 Effect of Weight Pruning across Different Layer and Module Types
All sub-figures clearly show an interesting phenomenon about these models in ICL inference: SVD-based weight pruning can enhance ICL performance in both shallow and deep layers across different module types. More surprising, deep layers weight matrices can be drastically reduced without much degradation in model performance, and may even get large improvements on it. This suggests that pruning weights in deep layers can effectively reduce model complexity while maintaining or enhancing model performance. And the model performance collapses to 0.5 (the expectation of a random guess in the binary task) after a sharp reduction at the shallow layers by contrast, indicating a higher sensitivity of the model to changes in shallow layers.
@@ 2.2 Effect of Different ICL Shot Numbers
Firstly, we note that without weight pruning, the performance of the model improves with an increase in the number of ICL shots. Which is consistent with prior works. Besides, for every shot number, a phenomenon consistent with what is described in Section 2.1 is observed: SVD-based weight pruning can enhance ICL performance, pruning deep layer weight is more stable than pruning shallow weight. Last but not the least, Figure 2 also demonstrates roughly that with a decrease in the number of shots, the rate of performance collapse in the model slows down after a sharp reduction at the shallow layer. Intuitively, this is because LLMs exhibit a shift in focus in the ICL setting, which results in a reduced scope of the output space. The more shots there are, the more pronounced the shift becomes, and this also leads to a faster collapse. We also theoretically discuss it in Section 3.2 and 3.3 (Remark 6).
@@ 3 Theoretical Analysis Results
In this section, we first describe the core components of our study by reviewing some basic notations in Section 3.1, and present the implicit gradient descent trajectories of ICL in Section 3.2. Afterwards, we give a mutual information based generalization bounds of ICL via full implicit GD trajectories in Section 3.3. Based on all our experimental and theoretical insights, we intuitively propose a derivative-free and effective method for downstream tasks in enhancing ICL inference in Section 3.4. Complete proofs can be found in the Appendix. For ease of qualitative analysis, our theoretical analysis is mainly focuses on linear attention setting. We discuss the standard Softmax attention setting in Appendix A.2 and feed-forward (MLP) layers in Appendix A.3.
@@ 3.1 Preliminaries
Expected generalization error. Given a loss function $\ell:\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}^{+}$, where $\ell(w,\mathbf{h})$ measures the "unfitness" or "error" of any $\mathbf{h}\in\mathcal{H}$ with respect to a hypothesis $w\in\mathcal{W}$. We take $\ell$ as a continuous function and assume that $\ell$ is differentiable almost everywhere with respect to $w$. The goal of learning is to find a hypothesis $w$ that minimizes the population risk, and for any $w\in\mathcal{W}$, the population risk is defined as $L_{\mu}(w)\triangleq\mathbb{E}_{\mathbf{h}\sim\mu}[\ell(w,\mathbf{h})].$ However, since $\mu$ via the sample $\mathbf{H}_{s}$ can only be partially observed, we instead turn to use the empirical risk, defined as $L_{\mathbf{H}_{s}}(w)\triangleq\frac{1}{N}\sum_{i=1}^{N}\ell(w,\mathbf{h}_{i})$. Then the expected generalization error of $\mathcal{A}$ is defined as
@@
In-context learning with Transformers [^2] Please refer to Appendix B.1 for an explanation on how Eq.(1) and Eq.(2) can utilize the same mask.. By prompt design, most context semantic understanding tasks can be unified into classification tasks. Simplify the form of each token in $\mathbf{H}$ to $\mathbf{h}_{i}=[\mathbf{x}_{i},\mathbf{y}_{i}]$, where $[\mathbf{x}_{i}]_{1}^{N}\in\mathbb{R}^{din}$ and $[\mathbf{y}_{i}]_{1}^{N}\in\mathbb{R}^{dout}$ are encoded input text and corresponding labels respectively. The test token has the form $\mathbf{h}_{N+1}=[\mathbf{x}_{N+1},mask]$, where $mask$ is the label needed to predict, and usually set 00 as its initialization. Therefore, the form of attention with residual connection333 In real-world scenarios, the residual connection module in Transformers is indispensable. is as follows:
@@ 3.2 The Implicit Gradient Descent Trajectories of ICL
To begin with, we present the implicit gradient descent of ICL, inspired by [21, 10, 35]. These works describe how ICL with attention can be connected to a meta-optimizer which produces implicit gradient. The following lemma demonstrates the result.
@@
Consider a Transformer consists of a single linear layer attention with residual connection, parameterized by $\mathbf{W}_{V}$,$\mathbf{W}_{K}$,$\mathbf{W}_{Q}$ as in Eq.(3). Same to Section 2, let $\mathbf{H}=[\mathbf{H}_{s},\mathbf{h}_{N+1}]$ be the input, where $\mathbf{H}_{s}=[\mathbf{h}_{1},\mathbf{h}_{2},...,\mathbf{h}_{N}]$ represents the demonstration sample and $\mathbf{h}_{N+1}$ is the test data. And let $\hat{\mathbf{h}}_{N+1}$ be the single layer output. Then, it holds that
@@
where $\mathbf{W}_{V}\mathbf{H}_{s}$ is regarded as the meta-gradient of ICL, which is used to generate the implicit gradient matrix $\Delta\mathbf{W}_{icl}$ to act on the final feature output. See Appendix A.1 for a proof.
@@
It is worth noting that $\text{rank}(\Delta\mathbf{W}_{icl})\leq\text{rank}(\mathbf{H}_{s})\leq N$ by rank relations in matrix multiplication, indicating that this is a low-rank operation. This explains why the effect of different ICL shot numbers is not uniform in Section 2.2. Details of the discussion in standard Softmax attention setting can be found in Appendix A.2. Plus, a significant body of work has discovered that the ICL capabilities of models are not robust to the order of ICL sample. For instance, Lu et al. [28] observed that large language models (LLMs) are sensitive to the sequence of ICL examples, and Liu et al. [27] reported that the context-based ICL performance exhibits a U-shaped curve—models tend to perform better with information that appears at the beginning or at the end of the input context. This aligns with observations from actual model training where the sequence of training samples affects outcomes especially when training with a batch size of $1$.
@@
where $\Delta\mathbf{W}_{t}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}^{t}{\mathbf{h}_{i}^{t-1}}\otimes\mathbf{W}_{K}^{t}{\mathbf{h}_{i}^{t-1}}\right)\mathbf{W}_{Q}^{t},$ $\mathbf{W}_{0}=0,\mathbf{W}_{t}=\mathbf{W}_{t-1}+\mathbf{G}_{t}$ and $\mathbf{G_{1}}=\Delta\mathbf{W}_{1}$. See Appendix A.4 for a proof.
@@ 3.3 Generalization Bounds of ICL via Full Implicit GD Trajectories
In addition, we also make a discussion on the noise of the implicit gradient in Section C.4, our analytical discussion indicates that the implicit gradients produced by Transformers in practical applications are noisy due to factors such as the extent of model pre-training and data characteristics (e.g., ICL shot number). We first present the assumption used in this subsection.
@@
where $\mathbf{vec(G}_{t})\in\mathbb{R}^{d}$, $\Delta\mathbf{W}_{t}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}^{t}{\mathbf{h}_{i}^{t-1}}\otimes\mathbf{W}_{K}^{t}{\mathbf{h}_{i}^{t-1}}\right)\mathbf{W}_{Q}^{t}$ and $\mathbf{G}_{t}=\Delta\mathbf{W}_{t}(1+\mathbf{W}_{0}+\sum_{j=1}^{t-1}\mathbf{G}_{j})=\Delta\mathbf{W}_{t}(1+\mathbf{W}_{t-1})$. $\mathrm{tr}\{\cdot\}$ denotes the trace of a matrix, $\left\lVert\cdot\right\rVert_{F}$ denotes the Frobenius norm of a matrix and $\mathbb{E}^{X}_{Y}$ is the conditional expectation. Proof details in Appendix A.5.
@@
Theorem 2 indicates that one can control the generalization performance of ICL via controlling the implicit gradient norm along the entire ICL implicit GD trajectories. Specifically, modulating the norm of $[\mathbf{G}_{t}]_{1}^{L}$ or $[\Delta\mathbf{W}_{t}]_{1}^{L}$ may enhance performance when utilizing ICL. Note that controlling implicit gradient norm can also control the magnitude of the trace of implicit gradient noise covariance. This elucidates why weight pruning through SVD, even if it only alters a single weight matrix, can confer advantages on the performance of Transformers in ICL inference. We will present an example below demonstrating how weight pruning can affect the norm of $\mathbf{G}_{t}$ or $\Delta\mathbf{W}_{t}$, thereby influencing the expected generalization error. Additional examples provided in Appendix A.6. This is also why, as illustrated in Figure 1, the highest model performance sometimes occurs at a clipping rate of 0.995. Furthermore, this could elucidate the utility of normalization in Transformers. However, weight pruning also impacts the expressive power, implying that increased pruning does not invariably lead to better outcomes. This clarifies why the highest model performance may also occur at clipping rates lower than 0.995.
@@
Consider $\Delta\mathbf{W}_{k}=\left(\sum_{i=1}^{N}\mathbf{W}_{V}^{k}{\mathbf{h}_{i}^{k-1}}\otimes\mathbf{W}_{K}^{k}{\mathbf{h}_{i}^{k-1}}\right)\mathbf{W}_{Q}^{k}$ when $t=k$. And we primarily consider the changes in a upper bound written as $UB(\left\lVert\Delta\mathbf{W}_{k}\right\rVert_{F}^{2})\triangleq\sum_{i=1}^{N}\left\lVert\mathbf{W}_{V}^{k}{\mathbf{h}_{i}^{k-1}}\otimes\mathbf{W}_{K}^{k}{\mathbf{h}_{i}^{k-1}}\right\rVert_{F}^{2}\left\lVert\mathbf{W}_{Q}^{k}\right\rVert_{F}^{2}$. Let $r$ represents the remained rank and $\delta$ represents the potential noise consisting of parts with small singular values, that is, $\mathbf{W}_{V}^{k}=\mathbf{W}_{V_{r}}^{k}+\delta_{V}=\mathbf{U}_{:r}^{V}\mathbf{\Sigma}_{:r}^{V}(\mathbf{V}_{:r}^{V})^{T}+\delta_{V}$, the same operation is applied to $\mathbf{W}_{Q}^{k}$ and $\mathbf{W}_{K}^{k}$ as well. Then we have
$\left\lVert\mathbf{W}_{V}^{k}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}$ $=\left\lVert(\mathbf{W}_{V_{r}}^{k}+\delta_{V})\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}$
$=\left\lVert\mathbf{W}_{V_{r}}^{k}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}+2(\mathbf{W}_{V_{r}}^{k}\mathbf{h}_{i}^{k-1})^{T}(\delta_{V}\mathbf{h}_{i}^{k-1})+\left\lVert\delta_{V}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}$
$=\left\lVert\mathbf{W}_{V_{r}}^{k}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}+\left\lVert\delta_{V}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2}\geq\left\lVert\mathbf{W}_{V_{r}}^{k}\mathbf{h}_{i}^{k-1}\right\rVert_{2}^{2},$
where $\mathbf{W}^{k}_{V_{r}}=\mathbf{U}_{:r}^{V}\mathbf{\Sigma}_{:r}^{V}(\mathbf{V}_{:r}^{V})^{T}$ and $\delta_{V}=\mathbf{U}_{r:}^{V}\mathbf{\Sigma}_{r:}^{V}(\mathbf{V}_{r:}^{V})^{T}$, and $\mathbf{U}_{:r},\mathbf{U}_{r:}$ are orthometric (properties of SVD), and further have
@@
where Eq. (*) is by [^5] For any vector $\mathbf{a}$ and $\mathbf{b}$, $\text{rank}(\mathbf{a}\otimes\mathbf{b})=1$, so the matrix ($\mathbf{a}\otimes\mathbf{b}$) only has one non-zero singular value. Combined with $||\mathbf{P}||_{F}=\sqrt{\sum_{i}\sigma_{i}^{2}(\mathbf{P})}$ and singular value is nonnegative, we can get $||\mathbf{a}\otimes\mathbf{b}||_{F}=\sqrt{\sum_{i}\sigma_{i}^{2}(\mathbf{a}\otimes\mathbf{b})}=\max_{i}[\sigma_{i}(\mathbf{a}\otimes\mathbf{b})]$, therefore, the unique non-zero singular value will decrease after performing SVD on $\mathbf{a}$ and/or $\mathbf{b}$.
@@
and Eq. (**) is by $\left\lVert\mathbf{P}\right\rVert_{F}=\sqrt{\sum_{i}\sigma_{i}^{2}(\mathbf{P})}$ for any matrix $\mathbf{A,B}$ and $\left\lVert\mathbf{P}\right\rVert_{F}\geq\left\lVert\mathbf{P}(r)\right\rVert_{F}$. $UB(\left\lVert\Delta\mathbf{W}_{k}(r)\right\rVert_{F}^{2})$ is the upper bound on $\lVert\Delta\mathbf{W}_{k}\lVert_{F}^{2}$ after using SVD.